# When the Correct Model Fails: The Optimality of Stackelberg Equilibria with Follower Intention Updates

### v1

@@ V-B1 Optimality against other beliefs
Figure 2: Percent of simulations with lowest cost achieved by each BR belief under OL information structure when $\tau=1$
@@ V-B2 Update time case study
Figure 3: Percentage of simulations where each BR belief obtains the lowest cost for $\tau=1,2,5,10,20$ under OL information structure
@@ V-C1 Optimality against other beliefs
This case study is identical to Figure 2 in setup, but the controls for both players come from the FSE solution (App. -A).
Figure 4: Percent of simulations with lowest cost achieved by BR beliefs under FB information structure and $\tau=1$
@@ V-C2 Estimation horizon case study
Figure 5: Percentage of simulations where each BR belief obtains the lowest cost for $\tau=1,2,5,10,20$ under FB information structure
@@ V-D Cost difference
Figure 6: Average percentage by which each belief’s cost exceeds the best model’s cost for each true BR and $\tau=1$ under OL and FB information structures
@@ V-E Discussion
This section discusses the results presented in Sections V-B and V-C. The first key observation is that Figures 2 and 4 show that the lowest cost is not always achieved by the true BR belief. Under both information structures, when the follower is either tracking or avoiding the leader, some alternative beliefs occasionally yield better performance.
@@ VI Conclusion
This paper examines the optimality conditions of Stackelberg OL and FB equilibria when the leader’s belief about the follower’s BR is updated over the game horizon. We prove sufficient conditions for the optimality of the true BR belief under the OL information structure and discuss its connection to the time-inconsistency of the OLSE. Under FB information structure, the SE is not necessarily Markov perfect, implying that when the leader holds incorrect BR beliefs, the resulting cost may, in some cases, be lower. Finally, we illustrate how these inconsistencies arise in a LQ game modeling a conflict-avoidance scenario. Future work will focus on characterizing the BR beliefs that may yield a lower cost than the true BR and identifying the conditions under which such outcomes occur.

### v2

@@ V-B1 Optimality against other beliefs
Figure 2: Percent of simulations with lowest cost achieved by each BR belief under OL information structure when $\tau=1$.
@@ V-B2 Update time case study
Figure 3: Percentage of simulations where each BR belief obtains the lowest cost for $\tau=1,2,5,10,20$ under OL information structure.
@@ V-C1 Optimality against other beliefs
This case study is identical to Figure 2 in setup, but the controls for both players come from the FSE solution (App. -A)
Figure 4: Percent of simulations with lowest cost achieved by BR beliefs under FB information structure and $\tau=1$.
@@ V-C2 Estimation horizon case study
Figure 5: Percentage of simulations where each BR belief obtains the lowest cost for $\tau=1,2,5,10,20$ under FB information structure.
@@ V-D Cost difference
Figure 6: Average percentage by which each belief’s cost exceeds the best model’s cost for each true BR and $\tau=1$ under OL and FB information structures.
@@ V-E Discussion
This section discusses the results presented in Sections V-B and V-C. The first key observation is that Figures 2 and 4 show that the lowest cost is not always achieved by the true BR belief, verifying our results from Sections IV-A and IV-B. Under both information structures, when the follower is either tracking or avoiding the leader, some alternative beliefs occasionally yield better performance.
@@ VI Conclusion
This paper examines the optimality conditions of Stackelberg OL and FB equilibria when the leader’s belief about the follower’s BR is updated over the game horizon. We prove sufficient conditions for the optimality of the true BR belief under the OL information structure and discuss its connection to the time-inconsistency of the OLSE. Under FB information structure, the SE is not necessarily Markov perfect, implying that when the leader holds incorrect BR beliefs, the resulting cost may, in some cases, be lower. Finally, we illustrate how these inconsistencies arise in a LQ game modeling a collision avoidance scenario. Future work will focus on characterizing the BR beliefs that may yield a lower cost than the true BR and identifying the conditions under which such outcomes occur.